Exercise 2: In a study, scientists examine data on the Sea temperature (variable r, in celsius) and Coral growth (responsPlease help with this statistics problem.
Answer #1

1. Graph A fits well the scatterplot

2. slope = change in y / change in x

Let us consider two points (26.1 , 9.2) and (26.3, 8.9)

slope, b = (8.9-9.2)/(26.3-26.1) = -1.5

y- intercept(a) is the point where the line meets the y axis

We know that y = a+bx

For a point (26.1 , 9.2)

9.2= a - 1.5*26.1

Solving , a= 48.35

3.Slope =-1.5

For each unit increase in x variable , y variable decreases by 1.5

4. the line of regression is

y=a+bx

growth = 48.35- 1.5*temperature

temp=30

growth = 48.35 - 30*1.5 = 3.35

5.We know that

For y =a +bx

b= r *(sy/sx)

using the values , b = -1.5 , r= -0.81 ,sx= 0.21

sy = 0.39

Also , a y-br

Thus , = 8.63